OpenStudy (anonymous):

I need help factoring 9x2+9x=4 and x4+3x2-28

OpenStudy (paxpolaris):

\[\Large 9x^2+9x=4 \]\[\implies \large 9x^2+9x-4 =0\]

OpenStudy (anonymous):

is that as far as you can go i set it up that way and got stuck

OpenStudy (paxpolaris):

you can split the middle term so that it adds to +9 and multiplies to -36

OpenStudy (paxpolaris):

or you can try using the cross method .... if you know it..

OpenStudy (anonymous):

cross method u really lost me lol sorry

OpenStudy (paxpolaris):

never mind... \[\implies \large 9x^2+12x-3x+4=0\]

OpenStudy (paxpolaris):

\[12-3=9 \]\[12 \times -3 = -36\]

OpenStudy (anonymous):